the length of a rectangle is 6 inches longer then twice the width. if the perimeter is 180 inches, find the demensions
Answer by CharlesG2(834)

the length of a rectangle is 6 inches longer then twice the width. if the perimeter is 180 inches, find the demensions
P perimeter = 2 * (L length + W width)
P = 2 * (2W + 6 + W) = 180
P = 2 * (3W + 6) = 180
2 * (3W + 6) = 180 (solve for W)
3W + 6 = 90
3W = 84
W = 28 inches
L = 2W + 6 = 2 * 28 + 6
L = 56 + 6
L = 62 inches
check:
28 * 2 + 62 * 2 = 56 + 124 = 180